40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Entry  Sat Jun 23 21:15:15 2018, Alena, General, Measurements, S1600594, S1600595, S1600597, S1600598 
    Reply  Sun Jun 24 14:47:50 2018, Gabriele, General, Measurements, S1600594, S1600595, S1600597, S1600598 
Message ID: 544     Entry time: Sun Jun 24 14:47:50 2018     In reply to: 543
Author: Gabriele 
Type: General 
Category: Measurements 
Subject: S1600594, S1600595, S1600597, S1600598 

Something was not quite right with the last measurement.

I wanted to check the QPD spectra (as explained in the procedure T1700089), but diaggui could not get the data, and failed with a "Timeout" error.

So I logged into the cymac and restarted the daqd process. In case you need to do that, in a terminal write

ssh cymac3 [ENTER]
killall daqd [ENTER]
exit [ENTER]

After this operation diaggui was able to read data and the QPD spectra looked ok. 

So I checked that enabling all fout ESD switches and eabling the ESD output did cause the disks to be excited.

Since everything looks normal, I started a new measurement, same parameters as before.

Quote:

Repeating the previous measurement after correcting the alignement. Did not vent the chamber.

  • 21:15 in chamber
    • S1600594 in CR1
    • S1600595in CR2
    • S1600597 in CR3
    • S1600598 in CR4

 

ELOG V3.1.3-